OpenStudy (anonymous):

is this correct ? cot^2x +1/cot^2x=csc^2x/cos^2x/sin^2x=1/sin^2x/cos^2x/sin^2x=cos^2x

OpenStudy (anonymous):

original question is cot^2x+1/cot^2x

OpenStudy (anonymous):

Using your trig identities: \[\cot^2 x+1=\csc^2 x\] \[\frac{\cot^2 x+1}{\cot^2 x}=\frac{\csc^2 x}{\cot^2 x}\] \[=\frac{1}{\sin^2 x} \times \frac{\sin^2 x }{\cos^2 x}\] \[=\frac{1}{\cos^2 x}\] \[=\csc^2 x\]

OpenStudy (anonymous):

so it reverst back to csc^2x?

OpenStudy (anonymous):

NO... \[\huge \frac{\csc^2}{\cot^2 x}=\frac{\frac{1}{\sin^2 x}}{\frac{1}{\tan^2 x}}\] Correct?

OpenStudy (anonymous):

but doesnt cotx = cosx/sinx?

OpenStudy (anonymous):

thats what Ihave in my notes..

OpenStudy (anonymous):

Separate into two fractions: \[\huge \frac{\frac{1}{\sin^2 x}}{\frac{1}{\tan^2 x}}=\frac{1}{\sin^2 x}\div \frac{1}{\tan^2 x}\] \[\huge =\frac{1}{\sin^2 x} \div \frac{\cos^2 x}{\sin^2 x}\] When you multiply you switch one fraction and get it's reciprocal. \[\huge =\frac{1}{\sin^2 x}\times \frac{\sin^2 x}{\cos^2 x}\]

OpenStudy (anonymous):

You didn't read your notes on the basics of changing a division sign into a multiplication sign it seems.

OpenStudy (anonymous):

which equal cos^2x

OpenStudy (anonymous):

How does: \[\frac{1}{\cos^2 x}\] equal \[\cos^2 x\]

OpenStudy (anonymous):

it doesn't thats just what my brain saw and computed

OpenStudy (anonymous):

So do you get it now that cos^2 x is not the answer?

OpenStudy (anonymous):

yeah I see where I went wrong.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

No worries.